# Balancing Act: Prioritization Strategies for LLM-Designed Restless Bandit Rewards

### v3

@@ Appendix A Additional Related Work
Multi-Objective Reinforcement Learning
Real-world decision-making often involves trade-offs between multiple, conflicting objectives, making multi-objective reinforcement learning (MORL) essential. A common approach in MORL is to scalarize multiple objectives into a single reward function, but this can oversimplify the problem. To address this, Pareto-based methods identify Pareto-optimal solutions that reflect different trade-offs between objectives, providing a more comprehensive decision-making framework Roijers et al. (2013); Van Moffaert and Nowé (2014). Recent advancements in MORL have introduced algorithms that enhance scalability, handle non-linear objectives, and improve exploration strategies. For example, Yang et al. (2019) proposed an algorithm that adapts policies based on changing objective importance. Additionally, empirical methods have been developed to evaluate MORL performance across scenarios Vamplew et al. (2011).
In multi-agent systems, MORL adds complexity as agents must balance their objectives while considering others’. Specialized approaches, such as extending Pareto-based methods to multi-agent contexts, aim to find Pareto-efficient solutions for all agents involved Bloembergen et al. (2015). Moreover, the concept of fairness becomes important in multi-agent MORL, as solutions must be equitable to all agents while still optimizing overall system performance. Researchers have begun to investigate fairness-aware MORL algorithms that explicitly consider the distribution of rewards among agents, ensuring that no single agent disproportionately benefits at the expense of others Jiang and Lu (2019); Zimmer et al. (2021).
(Pre-LLM) Reward Function Design
Designing effective reward functions for RL has long been a challenge due to the lack of robust learning signals. Ladosz et al. (2022). Previous approaches for reward design in RL include manually specifying reward components and tuning their coefficients Ng et al. (1999); Margolis and Agrawal (2023), and learning reward models from demonstrations Ho and Ermon (2016) or preferences Zhang and Ramponi (2023). While promising, these methods often demand substantial domain expertise, large amounts of data, or computationally expensive processes.
LLMs and Social Choice
In our adjudicator, we take a social choice perspective to model the problem of selecting a final reward function in a way that balances different objectives against each other. Principles from social choice have informed LLM-based algorithm in previous works, albeit in the very different context of creating a consensus statements for a group of users with diverse views Bakker et al. (2022); Fish et al. (2023). Moreover, social choice principles are also increasingly used to inform the fine-tuning of language models, including its alignment with human feedback Bakker et al. (2022); Conitzer et al. (2024); Ge et al. (2024); Dai and Fleisig (2024).
Appendix B Implementation Details
B.1. Whittle Index Policy
@@
To implement the Whittle Index computation, we use the method in Qian et al. (2016b) based on binary search. Additionally, for all the experiments, we cache computed whittle index for a given set of transition probabilities and reward functions to reduce computation time.
B.2. DLM Pipeline
In our work, we use a modified version of the DLM pipeline Behari et al. (2024), which employs the Whittle Index policy as the planner for RMAB. Our approach differs from Behari et al. (2024) in that we use the Whittle Index policy specifically for simulating RMAB, whereas Behari et al. (2024) use the PreFeRMAB policy (Zhao and Behari et al. 2023). This modification allows for faster and more stable simulations and effectively decouples the learning problem from the planning problem.
Specifically, the DLM consists of three components. First, a user provides a natural language preference $P$. We then create a prompt for LLM which includes the context of the RMAB problem, the preference $P$, a description of features available, and the index of those features in the feature array. Finally, the LLM is asked to generate the Python code of the reward function in text format. We describe the prompt used in Figures 6 and 7.
The LLM is queried $n_{p}$ times to obtain $n_{p}$ reward functions. For each reward function, we also compute the reward distribution over all the features. Next, given all the generated reward functions and their corresponding reward distributions, we query the LLM to select the best reward function. We describe the prompt used in Figures 8 and 9. This is called the reflection step. The best reward function is then used inside the prompt for next step of generating $n_{p}$ reward functions. This process is repeated $n_{r}$ times to obtain $n_{p}\cdot n_{r}$ reward functions. In all our reward function generation experiments, we query the LLM $n_{p}=4$ times and run the reflection loop $n_{r}=5$ times resulting in 20 candidate reward functions.
@@ B.2. DLM Pipeline
B.3. Synthetic Dataset Generation
@@
Table 2. Weight vector parameters for different synthetic datasets.
@@
B.4. Real World Dataset
As described in Section 6.1, one challenge in estimating transition probabilities under active actions is the rarity of such transitions due to the limited number of service calls. Consequently, the estimates for active transition probabilities tend to be noisy. To mitigate this issue, we follow a procedure similar to that in the synthetic domain to construct the datasets, but with some modifications. Specifically, instead of uniformly sampling passive probabilities in step 1, we use the transition probabilities estimated from real-world data Verma et al. (2023). Next, in step 2, we use the attributes describing the real-world beneficiaries to define features. Thus, instead of features A, B and C, we use features Age, Income and Education. Finally, we use the same weight vector values as in Table 2 to generate three real world datasets. This allows us to generate multiple datasets with varying levels of effect of intervention based on features while still using the realistic passive transition probabilities and feature distributions.
B.5. Hyperparameters
@@ B.6. Computing Resources
Appendix C Additional Results
C.1. Composite Prioritization Prompts
@@
Table 3. Results summary comparing different reward function choice strategies aggregated across the three synthetic Datasets. Cells in bold indicate the top 2 best values (higher is better). DLM: Decision-Language Model, SCLM: Social Choice Language Model, SCLM-SIM: Simulation based Scorer Model, SCLM-LLM: LLM based Scorer Model
@@
Table 4. Results summary comparing different reward function choice strategies aggregated across the three Real World Datasets. Cells in bold indicate the top 2 best values (higher is better). DLM: Decision-Language Model, SCLM: Social Choice Language Model, SCLM-SIM: Simulation based Scorer Model, SCLM-LLM: LLM based Scorer Model
@@
Table 5. Results comparing different reward function choice selection strategies for deciding tradeoff between preference alignment, unintended shift minimization and utility maximization, aggregated across the three synthetic datasets (left) and three real world datasets (right). A higher summed % change in desired features implies a better alignment, whereas less unintended shift is better.
@@
Table 6. Results comparing different reward function choice selection strategies for deciding tradeoff between preference alignment, unintended shift minimization and utility maximization, aggregated across the three synthetic datasets (left) and three real world datasets (right). A higher summed % change in desired features implies a better alignment, whereas less unintended shift is better.
@@
Appendix D Prompt Texts
In Figures 6-15, we show the prompts passed to LLM for various experiments described in the paper. Specifically, prompt in Figures 6,7 show how to generate a reward function in code form based on the problem description, the indices of features, and preference goals for synthetic and real-world problem domains, respectively. Figures 8, 9 show how the LLM is prompted to select the best reward function from those generated in the previous step for synthetic and real-world domains, respectively. Together, these prompts establish the DLM baseline in the paper. LLM-Zeroshot baseline has the exact same prompt as Figures 6, 7 for synthetic and real world domains. The only difference is that there is no reflection step, and the first reward generated by LLM is chosen as the best reward.
In Figure 10 and 11, we show how the prompt is enhanced with additional information to assist in selecting a reward function. The additional information is highlighted in bold. Section 6.4 explains how SCLM is used to specify additional objectives, such as minimizing utility shifts in features not included in the preference prompt and maximizing overall utility. In Figures 12, 13 we show the prompts for the DLM-EP baselines, which explicitly include instructions to minimize utility shifts in unintended features. In Figures 14, 15, we show the prompts that explicitly include instructions for maximizing overall utility.

### v4

@@ Appendix A Additional Related Work
Appendix B Additional Material for Section 4
(a) Income feature.
(b) Education feature.
Figure 5: Utility feature distributions for the default reward function (orange) and reward function returned for prompt “Prioritize agents with low income” (blue). The $x$-axis depicts the values of the feature and the $y$-axis the total utility generated by all agents with this feature value.
Unintended Utility Shifts
Moving from the default reward function to a reward function aligned with a given (prioritization) prompt causes shifts within the distribution of resources and utility. Due to correlations between features, this might change might lead to unintended utility shifts for features not specified in the prompt. In Figure 5, we show an instance of our experiments. We present the utility feature distribution for the two features income and education for two reward functions: The reward function selected by DLM for the prompt “ Prioritize agents with low income ” (orange) and the default reward (blue). While the utility generated by low-income agents increases, the utility generated by highly educated agents decreases, a side-effect a user is likely unaware of and that might conflict with their allocation goals.
Appendix C Implementation Details
C.1 Whittle Index Policy
@@
To implement the Whittle Index computation, we use the method in [26] based on binary search. Additionally, for all the experiments, we cache computed whittle index for a given set of transition probabilities and reward functions to reduce computation time.
C.2 DLM Pipeline
In our work, we use a modified version of the DLM pipeline [7], which employs the Whittle Index policy as the planner for RMAB. Our approach differs from Behari et al. [7] in that we use the Whittle Index policy specifically for simulating RMAB, whereas Behari et al. [7] use the PreFeRMAB policy (Zhao and Behari et al. 2023). This modification allows for faster and more stable simulations and effectively decouples the learning problem from the planning problem.
Specifically, the DLM consists of three components. First, a user provides a natural language preference $P$. We then create a prompt for LLM which includes the context of the RMAB problem, the preference $P$, a description of features available, and the index of those features in the feature array. Finally, the LLM is asked to generate the Python code of the reward function in text format. We describe the prompt used in Figures 7 and 8.
The LLM is queried $n_{p}$ times to obtain $n_{p}$ reward functions. For each reward function, we also compute the reward distribution over all the features. Next, given all the generated reward functions and their corresponding reward distributions, we query the LLM to select the best reward function. We describe the prompt used in Figures 9 and 10. This is called the reflection step. The best reward function is then used inside the prompt for next step of generating $n_{p}$ reward functions. This process is repeated $n_{r}$ times to obtain $n_{p}\cdot n_{r}$ reward functions. In all our reward function generation experiments, we query the LLM $n_{p}=4$ times and run the reflection loop $n_{r}=5$ times resulting in 20 candidate reward functions.
@@ C.2 DLM Pipeline
C.3 Synthetic Dataset Generation
@@
Table 2: Weight vector parameters for different synthetic datasets.
@@
C.4 Real World Dataset
As described in Section 6.1, one challenge in estimating transition probabilities under active actions is the rarity of such transitions due to the limited number of service calls. Consequently, the estimates for active transition probabilities tend to be noisy. To mitigate this issue, we follow a procedure similar to that in the synthetic domain to construct the datasets, but with some modifications. Specifically, instead of uniformly sampling passive probabilities in step 1, we use the transition probabilities estimated from real-world data [32]. Next, in step 2, we use the attributes describing the real-world beneficiaries to define features. Thus, instead of features A, B and C, we use features Age, Income and Education. Finally, we use the same weight vector values as in Table 2 to generate three real world datasets. This allows us to generate multiple datasets with varying levels of effect of intervention based on features while still using the realistic passive transition probabilities and feature distributions.
C.5 Hyperparameters
@@ C.6 Computing Resources
Appendix D Error Bound in Adjudicator
Multiplicative error
Suppose instead of observing the true score $s_{i}(R_{j})$, the Scorer Model (SCLM-SIM or SCLM-LLM) returns a noisy score estimate $\tilde{s}_{i}(R_{j})$ under multiplicative noise $\alpha_{i}^{j}\in(0,1]$ given as
$\alpha_{i}^{j}\cdot s_{i}(R_{j})\leq\tilde{s}_{i}(R_{j})\leq\frac{1}{\alpha_{i}^{j}}\cdot s_{i}(R_{j}),\quad\forall i\in[l],R_{j}\in\mathcal{R}.$ (6)
Let $\mathbf{s}(R_{j})=(s_{1}(R_{j}),s_{2}(R_{j}),...,s_{l}(R_{j}))$ represent the true score vector (and similarly, $\tilde{\mathbf{s}}(R_{j})$ the approximate score vector). If we consider $\alpha$ as the worst approximation factor across all scores, i.e., $\alpha=\min_{i\in[l],j\in[|\mathcal{R}|]}\alpha_{i}^{j}$, we can then write the vector form of the approximation as
$\alpha\cdot\mathbf{s}(R_{j})\leq\tilde{\mathbf{s}}(R_{j})\leq\frac{1}{\alpha}\cdot\mathbf{s}(R_{j}),\quad\forall R_{j}\in\mathcal{R}.$ (7)
This is true because $\alpha\cdot s_{i}(R_{j})\leq\alpha_{i}^{j}\cdot s_{i}(R_{j})$ and $1/\alpha_{i}^{j}\cdot s_{i}(R_{j})\leq 1/\alpha\cdot s_{i}(R_{j}),\;\forall i\in[l],R_{j}\in\mathcal{R}$.
Properties of Generalized p-mean
Lemma 2.
For all strictly positive vectors $\mathbf{x}\in\mathbb{R}^{N}_{>0}$, the generalized $p$-mean is:
  • positively homogeneous: $f_{p}(\lambda\mathbf{x})=\lambda f_{p}(\mathbf{x})$ for all $\lambda>0$, and
  • monotonic: if $\mathbf{x}\leq\mathbf{y}$ element wise, then $f_{p}(\mathbf{x})\leq f_{p}(\mathbf{y})$.
These properties hold for all $p\in[-\infty,1]$.
Proof.
Positive homogeneity:
Case 1: $p=-\infty$
In this case:
$f_{-\infty}(\mathbf{x})=\min_{i\in[N]}x_{i}$
Let $x_{j}=\min_{i\in[N]}x_{i}$. Then, for $\lambda>0$, it must be true that $\lambda x_{j}=\min_{i\in[N]}\lambda x_{i}$. Since $f_{-\infty}(\lambda\mathbf{x})=\lambda f_{-\infty}(\mathbf{x})$, $f_{-\infty}(\mathbf{x})$ is positive homogeneous.
Case 2: $p\not\in\{-\infty,0\}$
[…]
$f_{p}(\lambda\mathbf{x})=\left(\frac{1}{N}\sum_{i=1}^{N}(\lambda x_{i})^{p}\right)^{1/p}$ $=\left(\frac{1}{N}\sum_{i=1}^{N}\lambda^{p}x_{i}^{p}\right)^{1/p}$
$=\left(\lambda^{p}\cdot\frac{1}{N}\sum_{i=1}^{N}x_{i}^{p}\right)^{1/p}$
$=\lambda\left(\frac{1}{N}\sum_{i=1}^{N}x_{i}^{p}\right)^{1/p}=\lambda f_{p}(\mathbf{x})$
so $f_{p}$ is positively homogeneous.
Case 3: $p=0$
We use the definition of the geometric mean:
$f_{0}(\lambda\mathbf{x})=\left(\prod_{i=1}^{N}\lambda x_{i}\right)^{1/N}$ $=\left(\lambda^{N}\prod_{i=1}^{N}x_{i}\right)^{1/N}$
$=\lambda\left(\prod_{i=1}^{N}x_{i}\right)^{1/N}=\lambda f_{0}(\mathbf{x})$
Thus, positive homogeneity also holds for $p=0$.
Monotonicity: Suppose $\mathbf{x},\mathbf{y}\in\mathbb{R}^{N}_{>0}$ such that $x_{i}\leq y_{i}$ for all $i\in[N]$.
Case 1: $p=-\infty$
Let $x_{u}=\min_{i\in[N]}x_{i}$ and $y_{v}=\min_{i\in[N]}y_{i}$. Because $x_{u}$ is minimum over $\mathbf{x}$, $x_{u}\leq x_{v}$. Because $x_{i}\leq y_{i}$ for all $i\in[N]$, it must be true that $x_{v}\leq y_{v}$. Hence, $x_{u}\leq y_{v}$ or $\min_{i\in[N]}x_{i}\leq\min_{i\in[N]}y_{i}$ and thus $f_{-\infty}(\mathbf{x})$ is monotonic in $\mathbf{x}$.
Case 2: $p\notin\{-\infty,0\}$
Since the power function $u\mapsto u^{p}$ is monotonically increasing for all $p>0$ (and $u>0$), we have $x_{i}^{p}\leq y_{i}^{p}$ for all $i\in[N]$. Therefore,
$\frac{1}{N}\sum_{i=1}^{N}x_{i}^{p}\leq\frac{1}{N}\sum_{i=1}^{N}y_{i}^{p}$
Now, since $u\mapsto u^{1/p}$ is increasing for $p>0$, we can apply the power of $1/p$ and the inequality will be preserved because both sides are positive. Hence,
$\left(\frac{1}{N}\sum_{i\in[N]}x_{i}^{p}\right)^{\!1/p}\leq\left(\frac{1}{N}\sum_{i\in[N]}y_{i}^{p}\right)^{\!1/p}$
or
$f_{p}(\mathbf{x})\leq f_{p}(\mathbf{y})$
When $p<0$, $u\mapsto u^{p}$ is decreasing function (for u>0). Hence, we have $x_{i}^{p}\geq y_{i}^{p}$ for all $i\in[N]$.Therefore,
$\frac{1}{N}\sum_{i=1}^{N}x_{i}^{p}\geq\frac{1}{N}\sum_{i=1}^{N}y_{i}^{p}$
However, since $u\mapsto u^{1/p}$ is decreasing for $p<0$, we can apply the power of $1/p$ and the inequality will be flipped. Thus, we would get the same result as before
$\left(\frac{1}{N}\sum_{i\in[N]}x_{i}^{p}\right)^{\!1/p}\leq\left(\frac{1}{N}\sum_{i\in[N]}y_{i}^{p}\right)^{\!1/p}$
or
$f_{p}(\mathbf{x})\leq f_{p}(\mathbf{y})$
Case 3: $p=0$
We use the fact that that logarithm is a strictly increasing function (i.e., $u<v\Leftrightarrow\log u<\log v,\forall u,v>0$), and then reduce the product to sum
$\log f_{0}(\mathbf{x})$ $=\frac{1}{N}\sum_{i=1}^{N}\log x_{i}\leq\frac{1}{N}\sum_{i=1}^{N}\log y_{i}=\log f_{0}(\mathbf{y})$
$\Rightarrow\quad f_{0}(\mathbf{x})\leq f_{0}(\mathbf{y})$
Hence, the generalized p-mean function is both positively homogeneous and monotonic. ∎
Proof of Proposition 1:
Proof.
From our approximation definition in Inequality 2, we have
$\alpha\cdot\mathbf{s}(R_{j})\leq\tilde{\mathbf{s}}(R_{j})\leq\frac{1}{\alpha}\cdot\mathbf{s}(R_{j}),\quad\forall R_{j}\in\mathcal{R}.$
Because $f$ is monotonic (from Lemma 2), we can apply $f$ on both sides of the inequality without changing the direction. Thus we get
$f_{p}(\alpha\cdot\mathbf{s}(R_{j}))\leq f_{p}(\tilde{\mathbf{s}}(R_{j}))\leq f_{p}(\frac{1}{\alpha}\cdot\mathbf{s}(R_{j})),\quad\forall R_{j}\in\mathcal{R}.$ (8)
Next, because $f$ has positive homogeneity (from Lemma 2) for any scalar $\alpha>0$, we have:
$f_{p}(\alpha\cdot\mathbf{s}(R_{j}))=\alpha\cdot f_{p}(\mathbf{s}(R_{j})),$ (9)
and similarly,
$f_{p}\left(\frac{1}{\alpha}\cdot\mathbf{s}(R_{j})\right)=\frac{1}{\alpha}\cdot f_{p}(\mathbf{s}(R_{j})).$ (10)
Combining the above three observations, we get
$\alpha\cdot f_{p}(\mathbf{s}(R_{j}))\leq f_{p}(\tilde{\mathbf{s}}(R_{j}))\leq\frac{1}{\alpha}\cdot f_{p}(\mathbf{s}(R_{j})),\forall R_{j}\in\mathcal{R}.$ (11)
Since the Inequality 11 is true for all $R_{j}\in\mathcal{R}$, we can use $R_{j}=R^{*}$ to get:
$f_{p}(\tilde{\mathbf{s}}(R^{*}))\geq\alpha\cdot f_{p}(\mathbf{s}(R^{*})).$ (12)
Since $\tilde{R}^{*}$ maximizes $f_{p}(\tilde{\mathbf{s}}(R_{j}))$, we have:
$f_{p}(\tilde{\mathbf{s}}(\tilde{R}^{*}))\geq f_{p}(\tilde{\mathbf{s}}(R^{*})).$ (13)
Now, using the Inequality 11 for $R_{j}=\tilde{R}^{*}$, we get:
$f_{p}(\mathbf{s}(\tilde{R}^{*}))\geq\alpha\cdot f_{p}(\tilde{\mathbf{s}}(\tilde{R}^{*})).$ (14)
Combining the above three results, we get
$f_{p}(\mathbf{s}(\tilde{R}^{*}))\geq\alpha\cdot f_{p}(\tilde{\mathbf{s}}(\tilde{R}^{*}))\geq\alpha\cdot f_{p}(\tilde{\mathbf{s}}(R^{*}))\geq\alpha^{2}\cdot f_{p}(\mathbf{s}(R^{*})).$ (15)
Thus:
$\frac{f_{p}(\mathbf{s}(R^{*}))-f_{p}(\mathbf{s}(\tilde{R}^{*}))}{f_{p}(\mathbf{s}(R^{*}))}$ $\leq\frac{f_{p}(\mathbf{s}(R^{*}))-\alpha^{2}\cdot f_{p}(\mathbf{s}(R^{*}))}{f_{p}(\mathbf{s}(R^{*}))}$ (16)
$=\frac{(1-\alpha^{2})\cdot f_{p}(\mathbf{s}(R^{*}))}{f_{p}(\mathbf{s}(R^{*}))}=1-\alpha^{2}$ (17)
Relative $\textit{Relative Regret}=\frac{f_{p}(\mathbf{s}(R^{*}))-f_{p}(\mathbf{s}(\tilde{R}^{*}))}{f_{p}(\mathbf{s}(R^{*}))}\leq 1-\alpha^{2}$ (18)
Appendix E Additional Results
E.1 Composite Prioritization Prompts
@@
Table 3: Results summary comparing different reward function choice strategies aggregated across the three synthetic Datasets. Cells in bold indicate the top 2 best values (higher is better). DLM: Decision-Language Model, SCLM: Social Choice Language Model, SCLM-SIM: Simulation based Scorer Model, SCLM-LLM: LLM based Scorer Model
@@
Table 4: Results summary comparing different reward function choice strategies aggregated across the three Real World Datasets. Cells in bold indicate the top 2 best values (higher is better). DLM: Decision-Language Model, SCLM: Social Choice Language Model, SCLM-SIM: Simulation based Scorer Model, SCLM-LLM: LLM based Scorer Model
@@
Table 5: Results comparing different reward function choice selection strategies for deciding tradeoff between preference alignment, unintended shift minimization and utility maximization, aggregated across the three synthetic datasets (left) and three real world datasets (right). A higher summed % change in desired features implies a better alignment, whereas less unintended shift is better.
@@
Table 6: Results comparing different reward function choice selection strategies for deciding tradeoff between preference alignment, unintended shift minimization and utility maximization, aggregated across the three synthetic datasets (left) and three real world datasets (right). A higher summed % change in desired features implies a better alignment, whereas less unintended shift is better.
@@
E.2 Effect of Social Choice Function
Figure 6: Comparison of different methods using the prompt “Prioritize agents with a low value for feature $B$ and agents with a low value for feature $C$”.
While both the Utilitarian and Egalitarian social welfare function leads to good overall results, the choice between them significantly influences which reward function is selected at the instance level. Figure 6 shows one example instance for a prompt with two prioritization clauses. As in Figure 1, each point corresponds to a candidate reward function with the axes measuring alignment with the two respective clauses. (as described in the beginning of the section 6.3). The SCLM-Utilitarian model chooses a reward function from the Pareto frontier that shows a much stronger effect for the second clause (i.e., the utility increase for agents with a low value of feature $C$ is more pronounced). In contrast, the reward function selected by SCLM-Egalitarian is much more balanced. DLM-based baselines fail to pick a reward function from the Pareto frontier.
E.3 Preventing Ineffective Resource Use
Table 6 shows the results. Comparing DLM-PrioritizationOnly and DLM-ExtendedPrompt-Util, we again observe that adding the additional objective to the prompt does not result in a better performance in real-world domain and synthetic domain. In contrast, SCLM-ExtendedPrompt-Util which incorporates utility as a scoring dimension chooses reward functions resulting in a significantly higher utility increases for the prioritized subpopulations and a significantly higher total utility (i.e., there is a lower drop in utility).
Combining Prioritization, Unintended Shifts, and Utility Drops
Finally, we consider a variant of SCLM which both prevents unintended shifts and utility drops, i.e., in addition to the specified prioritization clauses, we add clause(s) on the minimization of shifts and a clause on the minimization of utility drops (which results in having a total of 4 evaluation dimensions; one per feature plus the utility drop one). We use the simulator scorer together with the utilitarian welfare function and call the resulting model SCLM-Full. We see in the last row of Table 6 that this model chooses balanced reward functions with high % change in utility in desired attributes, high overall utility, and low unintended shift, showcasing that SCLM can flexibly allow the end-user to combine multiple evaluation dimensions.
Appendix F Prompt Texts
In Figures 7-16, we show the prompts passed to LLM for various experiments described in the paper. Specifically, prompt in Figures 7,8 show how to generate a reward function in code form based on the problem description, the indices of features, and preference goals for synthetic and real-world problem domains, respectively. Figures 9, 10 show how the LLM is prompted to select the best reward function from those generated in the previous step for synthetic and real-world domains, respectively. Together, these prompts establish the DLM baseline in the paper. LLM-Zeroshot baseline has the exact same prompt as Figures 7, 8 for synthetic and real world domains. The only difference is that there is no reflection step, and the first reward generated by LLM is chosen as the best reward.
In Figure 11 and 12, we show how the prompt is enhanced with additional information to assist in selecting a reward function. The additional information is highlighted in bold. Section 6.4 explains how SCLM is used to specify additional objectives, such as minimizing utility shifts in features not included in the preference prompt and maximizing overall utility. In Figures 13, 14 we show the prompts for the DLM-EP baselines, which explicitly include instructions to minimize utility shifts in unintended features. In Figures 15, 16, we show the prompts that explicitly include instructions for maximizing overall utility.